# Number Cookbook: Number Understanding of Language Models and How to Improve It

### v2

@@ 3 How do tokenizers, PEs and data formats affect NUPA?
We have observed that the NUPA test poses significant challenges even for the most advanced LLMs. In this section, we aim to investigate the factors that can influence the NUPA of LLMs during their pretraining phase including tokenization strategies, PEs, and different data formats. We utilize the architecture of decoder-only transformers and alter the size to create models with 0.1B, 0.9B and 3B parameters. These models are trained from scratch, incorporating a wide range of techniques that could potentially impact NUPA. In this section, each model is trained on a single task . The details of the training process and models are included in Appendix A.4.1.
3.1 Tokenizer: One-digit tokenizers are good enough
Figure 3: Different tokenization of a long number. (a) mixed digit tokenizer, used in GPT-2. (b) one-digit tokenizer, used in Llama-2. (c) three-digit tokenizer, used in GPT-3.5, GPT-4 and Llama-3. (d) Aligned three-digit tokenizer.
LLMs interpret numbers as segmented tokens rather than whole numbers. With the development of language models, various tokenization strategies have emerged, including mixed tokenizers, one-digit tokenizers, and $k$-digit tokenizers ($k\geq 2$), as shown in Figure 3. In the BPE tokenizer used by GPT-2, the numbers are not specially treated, which resulted in irregular number cutting and is harmful to digit alignment. The cutting of numbers in modern tokenizers has become more aligned. These tokenizers greedily segment a number from left to right into $k$-digit tokens until a remainder shorter than $k$ digits is left, which is then segmented into a single token. Llama-2 use a one-digit tokenizer, but all of the latest LLMs use a tokenizer with $k=3$, subsequently coming with an extended vocabulary for numbers. Additonally, Singh & Strouse (2024) discover that just alternating the greedy direction from “left-to-right” to “right-to-left” (for integers) can improve performance of Llama-3 and GPT-4.
There is a growing tendency to expand the vocabulary size as the number of parameters in LLMs rapidly increases. Recent work has shown that a larger vocabulary is more suitable for larger LLMs (Tao et al., 2024) because longer tokens can encapsulate more complex and precise meanings for text tokens. But number is a different question:
@@ 3.1 Tokenizer: One-digit tokenizers are good enough
(a) 0.9B model int add
(b) 0.9B model float add
(c) 0.9B model int multiply
@@
From the figure, the one-digit tokenizer shows best in-domain performance in these three tasks, while three-digit tokenizer exhibits poor performance. In out-of-domain tests, one-digit tokenizer even exceeds the others by large margins. Tokenizers with an increasing number of digits significantly hinder sub-billion models’ NUPA. We also performed experiments on models of 3 different sizes including 0.1B, 0.9B, and 3B in Appendix A.4.2 and got similar results. Even as the model size increases, the performance of 2- or 3- digit tokenizer improves but remains either similar or worse than that of the one-digit tokenizers. For these experiments, we also report the digit match and dlength results in Appendix A.4.2 Figure 10 and 11, where one-digit tokenizer performs better both on digit learning (larger digit match) and length learning (less dlength). On the contrary, larger vocabularies significantly increase the model size requirements. In conclusion, we find no evidence to support the idea that increasing the vocabulary size improves NUPA performance.
@@
3.2 Special PEs are length regularization
Previous work has suggested that PE could be the key factor (Zhou et al., 2024) of length generalization. To further investigate whether the influence is specific on a certain task, we train 100M models with different PEs: RoPE (Su et al., 2023), NoPE (Kazemnejad et al., 2023a) and Alibi (Press et al., 2022) on four tasks: integer addition, float addition, fraction multiplication (easy) and scientific notation addition respectively. Models are trained on 1-8 lengths (S and M range), then test them on full range (S to XL, 1-20). RoPE, widely used in Llama and its derivatives, is the most classic relative PE. Then alibi, another relative PE, is proposed to address RoPE’s length overfitting issues. NoPE (transformers without PE, relying solely on the causal mask to encode the position information) offers a surprisingly easy way to achieve length generalization. Therefore, we compare these three typical PEs to evaluate the performance on NUPA.
@@ 3.3 Data formats help digit alignment
A series of works have proposed specific data formats including reverse formatting, zero padding and index hints. Reverse formatting (Lee et al., 2023; Shen et al., 2023) presents numbers in reverse order from the least significant digit to the most significant one to align with the models’ autoregressive mechanism, simplifying the learning process for addition. Zero padding (Lee et al., 2023; Shen et al., 2023; Zhou et al., 2024; Cho et al., 2024) adds leading zeros to numbers to standardize the lengths of operands, helping models align operands. Index Hints (Zhou et al., 2023) explicitly include positional information in input and output sequences by representing each digit along with its position index prefixing the digit.
While previous work mainly focuses on integer addition or multiplication, we extend the techniques to various tasks in NUPA Test of different number domains. To compare the effects of reverse formatting and zero padding, we demonstrate in Table 16 how the combination of reverse formatting and zero padding impacts length generalization. Reverse formatting, zero padding, and their combination all show better performance than vanilla formats in integer and float addition, while their performance is comparable to each other, suggesting that their functionality largely overlaps. Zero padding provides the function of helping alignment, and reverse formatting also plays a large role in alignment. The previously believed “helping calculation” function of reverse formatting is minor. As for index hint, we find it doesn’t work for our models. We discuss the details of these experiment results and the reasons in Appendix A.4.4.
@@ 3.4 Does finetuning improve NUPA performance of LLMs?
The existing techniques aimed at enhancing NUPA have rarely been applied to practical LLMs, mostly being tested on toy models and isolated tasks. This raises the question of whether it is possible to enhance the NUPA capabilities of large models through post-training finetuning. To explore this, we generate training sets ($10^{5}$ samples for each digit and each task) and validation sets for our NUPA tasks, ensuring no overlap with the original test set. We then used them to perform finetuning on a pre-trained model. Specifically, we finetune a Meta-Llama-3.1-8B model with lora (Hu et al., 2022) (rank 128, $\alpha$=32) on a mixed training set comprising all of our NUPA tasks. Remarkably, we find only 800 steps training (about 50M training samples, $\ll 1$ epoch) leads to significant improvement, as shown in Figure 2 with the finetuned model labeled as “Llama-8B-ft”. Though Llama-3.1-8B is not a strong baseline, this finetuned version achieves much better performance. For example, in max, max-hard, add-float and turediv tasks, this model even surpassed or matched GPT-4o, confirming our hypothesis: for many NUPA tasks, the model’s base capacity may not be the main limiting factor, but rather the lack of numerical diversity and task variety in the training data.
@@ 4 Is CoT suitable and valid for NUPA?
CoT has been proven to be effective to enhance the capacity of LLMs both theoretically (Feng et al., 2023; Yang et al., 2024b) and experimentally (Wei et al., 2023; OpenAI, 2024b). Thus, we are also interested in whether CoT is the ultimate solution for improving NUPA. Due to the task and representation diversity in our benchmark, it is hard to cover all issues with a single form of CoT. So we adapt a special CoT form called Rule-Following CoT (Hu et al., 2024) (RF-CoT), where LLMs are trained to follow a provided code or pseudo-code that outlines the procedure to solve the task. RF-CoT is capable of handling any problem with a solving procedure that can be broken down into recurrences and basic unit operations, making it well-suited for our benchmark tasks. The detailed introduction with an example of RF-CoT can be found in Appendix A.5.1.
@@
However, it requires a significantly longer context window and causes much slower inference speed compared to directly generating the answer. As shown in Table 3, with the 2000-token limit, CoT can only handle fraction addition involving numbers up to three digits. We provide the maximal digit length within the 2k context window limitation for each task in Appendix A.5.2 to show the context window limitation for complex tasks. As for inference time, Table 2 demonstrates the average inference time for generating each sample using “RF-CoT” and “direct answer” during the NUPA test where both experiments are operated on an A800 GPU. In the table, the “direct answer” with batch size $256$ shares a similar CUDA memory to RF-CoT with batch size $128$. The RF-CoT method is approximately 17 times slower than directly generating the answer, causing an unsustainable burden for such a basic operation that is frequently encountered in solving real-world problems, especially considering that number calculations may only account for a small part of a complex, real-world reasoning problem (such as analyzing a financial report).
@@ 5 Related work
Several works have probed numerical comprehension in encoder models. Wallace et al. (2019) trained probing models to to assess numerical understanding embedded in model representations, while Johnson et al. (2020) extend this conclusion to multi-language settings. Naik et al. (2019) used contrastive tests to evaluate models’ understanding of number magnitudes. Geva et al. (2020) demonstrated that fine-tuning on numerical reasoning data enhances the understanding. Unlike these studies, which focus on embeddings, our work emphasizes generating correct answers in autoregressive models. Recent efforts on such models include Razeghi et al. (2022), who studied few-shot learning correlations between term frequency and performance, and Zhang et al. (2024a), who identified key components in LLMs for basic arithmetic tasks. These works focus on some most classic tasks and our benchmark expands on these by incorporating diverse numerical representations, tasks, and digit ranges, offering a more comprehensive analysis.
@@
Datasets like those proposed by Spithourakis & Riedel (2018) and Lin et al. (2020) test numerical commonsense reasoning, while others focus on specific contexts, such as financial reasoning (Chen et al., 2022a, b) or tabular data (Akhtar et al., 2023). These works highlight numerical reasoning within specific domains rather than general numerical processing tasks. In contrast, our benchmark targets core numerical understanding, emphasizing tasks decoupled from domain-specific constraints.
@@
Despite its close relationship with NUPA, mathematical reasoning is a broader field involving diverse skills such as task comprehension, equation solving, tool usage, and more (Lu et al., 2023b). While correct numerical processing is a critical component of mathematical reasoning, it is not the entirety of it (Stolfo et al., 2023). Datasets like MathQA (Amini et al., 2019), GSM8k (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and SVAMP (Patel et al., 2021) focus on math word problems requiring multi-step reasoning and problem-solving. Few works isolate numerical processing from mathematical reasoning. Saxton et al. (2019) introduced a dataset for numerical tasks, such as adding floating-point numbers, but lacked task categorization by difficulty or length. Moreover, mixing numerical and algebraic tasks complicated analyses of pure numerical processing. Our benchmark addresses this gap, offering fine-grained categorization and evaluation of numerical understanding tasks.
@@ 6 Conclusion
We investigate NUPA of LLMs and introduce a comprehensive benchmark, the NUPA test, to reveal that numerical problems remain challenging for modern LLMs. Our comprehensive test suite, which includes a variety of numerical representations and tasks, has exposed the surprising vulnerability of LLMs in this fundamental area. To explore ways to improve NUPA, we extend and evaluate previous pre-training techniques on the NUPA benchmark. While direct finetuning on the NUPA tasks does improve the performance, using those tricks specifically designed for NUPA in the finetuning tends to harm it, suggesting that these methods are not easily transferable to practical LLMs. We also explore the potential of chain-of-thought techniques to enhance NUPA and discuss their limitations.
@@ 7 Limitation
As a benchmark that specifically focuses on number understanding and processing abilities, we acknowledge that the range of tasks could still be incomplete and biased toward certain aspects. We will continue updating our benchmark, including but not limited to adding new tasks and refining existing ones to ensure appropriate difficulty. Additionally, the number of models we have tested so far is limited, and we plan to include more promising pre-trained models in future evaluations.
On the other hand, although we have identified the limitations of LLMs’ NUPA, the existing solutions each have their own drawbacks. We have yet to find a path that fully addresses the problem. Solving this issue may require research across multiple fields, such as enhancing the diversity of pre-training corpora, developing new techniques, or enabling more efficient reasoning paradigms that make more complex CoT approaches feasible. We hope our work can contribute to and be complemented by advancements in these areas.
@@ A.1.1 Representations
  • Fraction: A fraction has two parts: numerator and denominator and we use a “/” to separate the integer and decimal parts. Unless otherwise specified, all fractions mentioned in this paper are in their simplest form (that is the numerator and denominator are coprime), but they may be greater than 1. An example is 12/7. Only in the “truediv” task between two fractions, because the “/” is also the division operator, We enclose fractions in a pair of parentheses like (12/7) / (2/3) = 18/7 to make it clear.
  • Scientific Notation: A scientific notation has two parts: significand and exponent. In our benchmark, the significand is always a float larger than 1 and less than 10 and the exponent should be a positive integer (and we also set an upper bound of 99). We use a “e” to seperate these two parts. An example is 1.5e8.

### v3

@@ 3 How do tokenizers, PEs and data formats affect NUPA?
We have observed that the NUPA Test poses significant challenges even for the most advanced LLMs. In this section, we aim to investigate the factors that can influence the NUPA of LLMs during their pretraining phase, including tokenization strategies, PEs, and different data formats. We utilize the architecture of decoder-only transformers and alter the size to create models with 0.1B, 0.9B and 3B parameters. These models are trained from scratch, incorporating a wide range of techniques that could potentially impact NUPA. In this section, each model is trained on a single task . The details of the training process and models are included in Appendix A.4.1.
3.1 One-digit tokenizers are good enough
Figure 3: Different tokenization of a long number. (a) GPT2: mixed digit tokenizer, (b) Llama-2: one-digit tokenizer. (c) GPT-3.5, GPT-4 and Llama-3: three-digit tokenizer. (d) Aligned three-digit tokenizer.
LLMs interpret numbers as segmented tokens rather than whole numbers. With the development of language models, various tokenization strategies have emerged, including mixed tokenizers, one-digit tokenizers, and $k$-digit tokenizers ($k\geq 2$), as shown in Figure 3. In the BPE tokenizer used by GPT-2, the numbers are not specially treated, which resulted in irregular number cutting and is harmful to digit alignment. The cutting of numbers in modern tokenizers has become more aligned. These tokenizers greedily segment a number from left to right into $k$-digit tokens until a remainder shorter than $k$ digits is left, which is then segmented into a single token. Llama-2 uses a one-digit tokenizer, but all of the latest LLMs use a tokenizer with $k=3$, which comes with an extended vocabulary for numbers. Additionally, Singh & Strouse (2024) discovers that just alternating the greedy direction from “left-to-right” to “right-to-left” (for integers) can improve performance of Llama-3 and GPT-4.
There is a growing tendency to expand the vocabulary size as the number of parameters in LLMs rapidly increases. Recent work has shown that a larger vocabulary is more suitable for larger LLMs (Tao et al., 2025) because longer tokens can encapsulate more complex and precise meanings for text tokens. However, numbers behave differently:
@@ 3.1 One-digit tokenizers are good enough
(a) int add
(b) float add
(c) int multiply
@@
From the figure, the one-digit tokenizer shows the best in-domain performance in these three tasks, while three-digit tokenizer exhibits poor performance. In out-of-domain tests, the one-digit tokenizer also exceeds the others by large margins. Tokenizers with an increasing number of digits significantly hinder sub-billion models’ NUPA. We also performed experiments on models of 3 different sizes including 0.1B, 0.9B, and 3B in Appendix A.4.2 and got similar results. Even as the model size increases, the performance of 2- or 3- digit tokenizer improves but remains either similar or worse than that of the one-digit tokenizers. For these experiments, we also report the digit match and dlength results in Appendix A.4.2 Figure 10 and 11, where one-digit tokenizer performs better both on digit learning (larger digit match) and length learning (less dlength). On the contrary, larger vocabularies significantly increase the model size requirements. In conclusion, we find no evidence to support the idea that increasing the vocabulary size improves NUPA performance.
@@
3.2 Specialized PEs Act as Length Regularizers
Previous work has suggested that PE could be the key factor (Zhou et al., 2024b) of length generalization. To further investigate whether the influence is specific on a certain task, we train 100M models with different PEs: RoPE (Su et al., 2024), NoPE (Kazemnejad et al., 2023) and Alibi (Press et al., 2022) on four tasks: integer addition, float addition, fraction multiplication (easy) and scientific notation addition respectively. Models are trained on 1-8 lengths (S and M range), then test them on full range (S to XL, 1-20). RoPE, widely used in Llama and its derivatives, is the most classic relative PE. Then Alibi, another relative PE, is proposed to address RoPE’s length overfitting issues. NoPE (transformers without PE, relying solely on the causal mask to encode the position information) offers a surprisingly easy way to achieve length generalization. Therefore, we compare these three typical PEs to evaluate the performance on NUPA.
@@ 3.3 Data formats help digit alignment
A series of works have proposed specific data formats including reverse formatting, zero padding and index hints. Reverse formatting (Lee et al., 2024; Shen et al., 2023) presents numbers in reverse order from the least significant digit to the most significant one to align with the models’ autoregressive mechanism, simplifying the learning process for addition. Zero padding (Lee et al., 2024; Shen et al., 2023; Zhou et al., 2024b; Cho et al., 2024) adds leading zeros to numbers to standardize the lengths of operands, helping models align operands. Index Hints (Zhou et al., 2024a) explicitly incorporate positional information by prefixing each digit with its corresponding position index in both input and output sequences.
While previous work mainly focuses on integer addition or multiplication, we extend the techniques to various tasks in the NUPA Test of different number domains. To compare the effects of reverse formatting and zero padding, we demonstrate in Table 16 how the combination of reverse formatting and zero padding impacts length generalization. Reverse formatting, zero padding, and their combination all outperform vanilla formats in integer and float addition, while their performance is comparable to each other, suggesting that their functionality largely overlaps. Zero padding helps ensure proper alignment, while reverse formatting also plays a crucial role in maintaining alignment. The previously believed “helping calculation” function of reverse formatting is minor. As for index hint, we find it doesn’t work for our models. We discuss the details of these experiment results and the reasons in Appendix A.4.4.
@@ 3.4 Does finetuning improve NUPA performance of LLMs?
The existing techniques aimed at enhancing NUPA have rarely been applied to practical LLMs, mostly being tested on toy models and isolated tasks. This raises the question of whether it is possible to enhance the NUPA capabilities of large models through post-training finetuning. To explore this, we generate training sets ($10^{5}$ samples for each digit and each task) and validation sets for our NUPA tasks, ensuring no overlap with the original test set. We then use them to finetune a pretrained model. Specifically, we finetune a Meta-Llama-3.1-8B model with LoRA (Hu et al., 2022) (rank 128, $\alpha$=32) on a mixed training set comprising all of our NUPA tasks. Remarkably, we find only 800 steps training (about 50M training samples, $\ll 1$ epoch) leads to significant improvement, as shown in Figure 2 with the finetuned model labeled as “Llama-8B-ft”. Though Llama-3.1-8B is not a strong baseline, this finetuned version achieves much better performance. For example, in max, max-hard, add-float and truediv tasks, this model even surpassed or matched GPT-4o, confirming our hypothesis: for many NUPA tasks, the model’s base capacity may not be the main limiting factor, but rather the lack of numerical diversity and task variety in the training data.
@@ 4 Is CoT suitable and valid for NUPA?
CoT has been proven to be effective in enhancing the capacity of LLMs both theoretically (Feng et al., 2023; Yang et al., 2024b) and experimentally (Wei et al., 2022; OpenAI, 2024b). Thus, we are also interested in whether CoT is the ultimate solution for improving NUPA. Due to the task and representation diversity in our benchmark, it is hard to cover all issues with a single form of CoT. So we adapt a special CoT form called Rule-Following CoT (Hu et al., 2024) (RF-CoT), where LLMs are trained to follow a provided code or pseudo-code that outlines the procedure to solve the task. RF-CoT is capable of handling any problem with a solving procedure that can be broken down into recurrences and basic unit operations, making it well-suited for our benchmark tasks. The detailed introduction with an example of RF-CoT can be found in Appendix A.5.1.
@@
However, it requires a significantly longer context window and causes much slower inference speed compared to directly generating the answer. As shown in Table 3, with the 2000-token limit, CoT can only handle fraction addition involving numbers up to three digits. We provide the maximal digit length within the 2k context window limitation for each task in Appendix A.5.2 to show the context window limitation for complex tasks. As for inference time, Table 2 demonstrates the average inference time for generating each sample using “RF-CoT” and “direct answer” during the NUPA Test where both experiments are operated on an A800 GPU. In the table, the “direct answer” with batch size $256$ uses a similar amount of CUDA memory as RF-CoT with batch size $128$. The RF-CoT method is approximately 17 times slower than directly generating the answer, causing an unsustainable burden for such a basic operation that is frequently encountered in solving real-world problems, especially considering that number calculations may only account for a small part of a complex, real-world reasoning problem (such as analyzing a financial report).
@@ 5 Related work
Several works have probed numerical comprehension in encoder models. Wallace et al. (2019) trained probing models to assess numerical understanding embedded in model representations, while Johnson et al. (2020) extend this conclusion to multi-language settings. Naik et al. (2019) used contrastive tests to evaluate models’ understanding of number magnitudes. Geva et al. (2020) demonstrated that finetuning on numerical reasoning data enhances the understanding. Unlike these studies, which focus on embeddings, our work emphasizes generating correct answers in autoregressive models. Recent efforts on such models include Razeghi et al. (2022), who studied few-shot learning correlations between term frequency and performance, and Zhang et al. (2024a), who identified key components in LLMs for basic arithmetic tasks. These works focus on some most classic tasks and our benchmark expands on these by incorporating diverse numerical representations, tasks, and digit ranges, offering a more comprehensive analysis.
@@
Datasets like those proposed by Spithourakis & Riedel (2018) and Lin et al. (2020) test numerical commonsense reasoning, while others focus on specific contexts, such as financial reasoning (Chen et al., 2021; 2022) or tabular data (Akhtar et al., 2023). These works highlight numerical reasoning within specific domains rather than general numerical processing tasks. In contrast, our benchmark targets core numerical understanding, emphasizing tasks decoupled from domain-specific constraints.
@@
Despite its close relationship with NUPA, mathematical reasoning is a broader field involving diverse skills such as task comprehension, equation solving, tool usage, and more (Lu et al., 2023b). While correct numerical processing is a critical component of mathematical reasoning, it is not the entirety of it (Stolfo et al., 2023). Datasets like MathQA (Amini et al., 2019), GSM8k (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), and SVAMP (Patel et al., 2021) focus on math word problems requiring multi-step reasoning and problem-solving. Few works isolate numerical processing from mathematical reasoning. Saxton et al. (2019) introduced a dataset for numerical tasks, such as adding floating-point numbers, but lacked task categorization by difficulty or length. Moreover, mixing numerical and algebraic tasks complicated analyses of pure numerical processing. Our benchmark addresses this gap, offering fine-grained categorization and evaluation of numerical understanding tasks.
@@ 6 Conclusion
We investigate NUPA of LLMs and introduce a comprehensive benchmark, the NUPA Test, to reveal that numerical problems remain challenging for modern LLMs. Our comprehensive test, which includes a variety of numerical representations and tasks, has exposed the surprising vulnerability of LLMs in this fundamental area. To explore ways to improve NUPA, we extend and evaluate previous pretraining techniques on the NUPA benchmark. While direct finetuning on the NUPA tasks does improve the performance, utilizing those tricks specifically designed for NUPA in the finetuning tends to harm NUPA, suggesting that these methods are not easily transferable to practical LLMs. We also explore the potential of chain-of-thought techniques to enhance NUPA and discuss their limitations.
@@ 7 Limitation
As a benchmark that specifically focuses on number understanding and processing abilities, we acknowledge that the range of tasks could still be incomplete and biased toward certain aspects. We will continue updating our benchmark, including but not limited to adding new tasks and refining existing ones to ensure appropriate difficulty. Additionally, the number of models we have tested so far is limited, and we plan to include more promising pretrained models in future evaluations.
On the other hand, although we have identified the limitations of LLMs’ NUPA, the existing solutions each have their own drawbacks. We have yet to find a path that fully addresses the problem. Solving this issue may require research across multiple fields, such as enhancing the diversity of pretraining corpora, developing new techniques, or enabling more efficient reasoning paradigms that make more complex CoT approaches feasible. We hope our work can contribute to and be complemented by advancements in these areas.
Reproducibility Statement
We have made every effort to ensure that the results presented in this paper are fully reproducible. Detailed descriptions of the number formats, construction and metrics of our NUPA dataset are provided in Section 2 and A.1.5, and examples for each task in A.4.4. To further facilitate reproducibility, we have included the complete dataset and the source code, enabling the generation of the entire dataset and the training and assessment of models, within the supplementary materials and the github page https://github.com/GraphPKU/number_cookbook. Researchers wishing to generate NUPA benchmark or replicate our experiments can refer to these resources for all necessary information.
Ethics Statement
In conducting this research, we have adhered to the highest ethical standards to ensure the integrity and fairness of our work. For source code releases, we have ensured compliance with applicable legal standards. During the construction of the dataset, all data was entirely generated randomly, without including any personal identity information or other private data of individuals.
Acknowledgements
This work was supported by National Key R&D Program of China (2022ZD0160300) and the NSF China (No. 62276004).
@@ A.1.1 Representations
  • Fraction: A fraction has two parts: numerator and denominator and we use a “/” to separate the numerator and denominator parts. Unless otherwise specified, all fractions mentioned in this paper are in their simplest form (that is the numerator and denominator are coprime), but they may be greater than 1. An example is 12/7. Only in the “truediv” task between two fractions, because the “/” is also the division operator, we enclose fractions in a pair of parentheses like (12/7) / (2/3) = 18/7 to make it clear.
  • Scientific Notation: A scientific notation has two parts: significand and exponent. In our benchmark, the significand is always a float larger than 1 and less than 10 and the exponent should be a positive integer (and we also set an upper bound of 99). We use a “e” to separate these two parts. An example is 1.5e8.